What is the equivalent of P (6 ≤ X ≤ 8)?


A) P(X ≤ 8) - P(X ≤ 6)

B) P(X ≤ 8) - P(X ≤ 5)

C) 1- [P(X ≤ 8) + P(X ≤ 6)]

D) P(X ≤ 8) + P(X ≤ 6)


Choose The Right answer


1
Expert's answer
2021-05-07T13:42:02-0400

"P(6\leq X\leq8)=P(\dfrac{6-8}{2}\leq Z<\dfrac{8-8}{2})=P(-1\leq Z<0)=0.5-0.1587=0.3423"




"P(X\leq 8)=P(Z<\dfrac{8-8}{2})=P(Z<0)=0.5\\P(X\leq 6)=P(Z<\dfrac{6-8}{2})=P(Z<-1)=0.1587"


So, "P(6\leq X\leq 8)=P(X\leq 8)-P(X\leq 6)"


Answer: Option A is correct